# 1]organization=Scuola Normale Superiore and INFN, addressline=Piazza dei Cavalieri 7, city=Pisa, postcode=56126, state=PI, country=Italy

### v3

@@ Abstract
Introduction.—Assuming that the early universe underwent a phase of cosmic inflation, fields that weakly couple to the inflaton can play a crucial role in shaping the primordial non-Gaussianities—also referred to as cosmological correlators. One possible origin of scalar perturbations during inflation is the presence of composite particles, which arise from strongly coupled dynamics among more elementary constituents, similar to the mechanisms in quantum chromodynamics (QCD) and models of the composite Higgs. This scenario has been far less explored compared to the more conventional treatments of cosmological correlators involving fundamental particle exchanges. For recent developments in this direction, see, for example, [1, 2, 3, 4, 5].
@@
Setup.—To investigate a strongly coupled DS during inflation, it is useful to classify different scenarios according to the mass gap: for example, $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$, $\text{M}_{\text{Gap}}\sim H_{\text{inf}}$ and $\text{M}_{\text{Gap}}\gg H_{\text{inf}}$. In this work, the Dark Walker model falls into the $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$ regime, where the DS is well approximated by a conformal field theory. Such a conformal sector may contain operators with large anomalous dimensions, which are called unparticles, as introduced in [6, 7], etc. An extended notion of unparticles with a mass gap has also been discussed in [8], but lies beyond the scope of this work. Gapless unparticles, however, cannot serve as DM, since their behavior resembles radiation rather than particles. Thus, the Dark Walker has to experience a phase transition, forming gapped states. During inflation, the model remains near the infrared fixed point (IRFP) long enough to leave imprints in the primordial non-Gaussianity. This behavior exemplifies a form of “walking dynamics”, which gives rise to the name “Dark Walker”.
@@
It is long known that there exists a non-trivial IRFP for $\text{SU}(N_{c})$ theories with a certain range of $N_{f}$, which is called the conformal window [9]. The $\beta$-function of $\text{SU}(N_{c})$ theories in the asymptotically free phase is given in [10, 11]
@@
From the $\beta$-function we notice that, generally in the perturbative regime $g\lesssim 4\pi$, the IRFP is stable [9], which means that a small relevant deformation will not drag the Dark Walker into the gapped phase. Therefore, we have to look for special gauge theories. Fortunately, recent studies in lattice QCD (e.g. [12, 13]) have revealed that, for the SU(2) theory with four flavors of massless Dirac fermions and the $\text{SU}(3)$ theory with eight flavors of massless Dirac fermions, the ultraviolet (UV) and IR fixed points will merge, as illustrated in Figure 1.
@@
In the following, we will take the Dark Walker to be an $\text{SU}(3)_{D}$ theory with $N_{f}=8$ flavors of massless Dirac fermions. This is an illustrator of our Dark Walker mechanism. In general, this scenario should hold for all gauge models with a merged fixed point (MFP). This kind of mass generation is called the symmetric mass generation (SMG), a phenomenon where the fermions acquire mass gaps while the chiral symmetry remains unbroken. For a recent review, the readers can refer to [14].
During Inflation.—In this work, we choose the value of $H_{\text{inf}}$ to be the current upper bound value from [15], which is $H_{\text{inf}}\sim 2.5\times 10^{-5}M_{\text{pl}}\sim 6.0\times 10^{13}\text{ GeV}$. In the context of cosmology, it is convenient to use the number of $e$-folds NN to measure the ratio between scales. There are $\sim 10$ $e$-folds between the Planck mass scale and the Hubble scale during inflation. Once we assume the inflation starts from the $M_{\text{pl}}$, the working energy scale during inflation flows from $M_{\text{pl}}$ to $H_{\text{inf}}$, and stays at $H_{\text{inf}}$ until the end of inflation.
@@
We will take the critical gauge coupling $g_{c}^{2}$ to be the lattice result of [13]. The critical gradient-flow coupling $g_{\text{GF}}^{2}\approx 25$ will give us $g_{c}^{2}\lesssim 12$, with reference to [16]. Surely, this one-loop matching in [16] will break up for strong couplings, but all we need is an upper bound for the number of $e$-folds needed. We will restrict the ranges of $N_{\text{cw}}$ and the initial value of the gauge coupling $g_{0}$ after constructing the effective interactions.
Once our model reaches the MFP, the gauge theory (1) becomes a conformal theory, and we cannot use this Lagrangian. Generally, it is not easy to derive the operator spectrum and the operator product expansion (OPE) coefficients between operators from the UV theory (1). Nevertheless, we can focus on a few operators that we expect to see from the UV theory. An important one is the fermion bilinear operator $\mathcal{O}_{\psi}\equiv\bar{\psi}\psi$, as the anomalous dimension $\gamma_{m}$ of $\mathcal{O}_{\psi}$ will be greater than 1 near the MFP, studied in [17]. This property will trigger “walking” and drag the Dark Walker into the gapped phase, as we will see below. Another possible primary scalar operator is $\mathcal{O}_{F}\equiv\text{Tr}F^{2}$, where the scaling dimension $\Delta_{F}\lesssim 4$. The explicit value can only be found from lattice simulations. The operator $\mathcal{O}_{\psi}$ has the scaling dimension $\Delta_{\psi}=3-\gamma_{m}\approx 1.8$ from [17], and we assume $\Delta_{F}\approx 3.9$ for $\mathcal{O}_{F}$.
Using the language of effective field theory (EFT) of inflation [18, 19], we can construct the linear interactions between the scalar primary operators $\mathcal{O}_{\Delta}$ and the scalar perturbation $\pi$ as given in [1, 5]:
@@
In addition, assuming that $\langle\mathcal{O}_{\Delta}^{3}\rangle$ is negligible, the bispectrum is determined by the quadratic derivative coupling $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\Delta}$, as discussed in [5]. Having established the EFT of inflation with unparticles, we now turn to its cosmological signatures.
Non-Gaussianity.—The cosmological bispectrum can be calculated using the results in [5]. The non-Gaussianity $f_{\text{NL}}$ is defined as
@@
is the three-point correlation function of curvature perturbation $\zeta$. $\Delta_{\zeta}^{2}=2.1\times 10^{-9}$ is the dimensionless power spectrum of $\zeta$ from [20]. At linear order, the curvature perturbation $\zeta$ relates to the scalar perturbation $\pi$ via
@@
Taking the slow-roll parameter $\epsilon$ to be $0.0097$ according to [15], we find that $1.2\times 10^{-8}\lesssim f_{\text{NL},\psi}\lesssim-13$ and $|f_{\text{NL},F}|<10^{-10}$. The full expression of $f_{\text{NL}}$ is presented in Appendix A. The $f_{\text{NL}}$ with $\mathcal{O}_{\psi}$ is desirable because it has the potential to be $O(1)$, which is around the center of the latest observation $f_{\text{NL}}=6\pm 46$ [21], while large enough to be detectable with the experiments in the near future. The $f_{\text{NL},F}$ is too small to be detected. Even the futuristic 21-cm tomography can only give a limiting strength of non-Gaussianity for observability about $f_{\text{NL}}\gtrsim 10^{-4}$ [22]. Notice that the choice of $H_{\text{inf}}$ will significantly affect the value of $f_{\text{NL}}$, as indicated in Figure 2.
@@
Figure 3 shows the shape functions with $\mathcal{O}_{\psi}$ and $\mathcal{O}_{F}$ exchanges. Both of them are near-equilateral, while the squeezed limit behaviors degenerate with the equilateral non-Gaussianity. Only the full shapes of the bispectra can break this degeneracy, as emphasized in [5].
@@
After inflation.—This section connects the inflationary phase to the present-day dark matter abundance. After the inflation ends, the universe becomes almost flat. The interactions $\dot{\pi}\mathcal{O}_{\psi}$ and $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\psi}$ will freeze, since the energy scale becomes too small compared to the EFT of inflation scale $M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}$. However, even though the linear mixing interaction $\dot{\pi}\mathcal{O}_{\psi}$ remains small before it freezes, its nature as a relevant operator will drag our model into the gapped phase. The gapped states are natural DM candidates. The masses of the lowest lying states of pseudoscalars (PSs) and vector mesons (Vs) can be read from the lattice QCD results [13, 23], where the mass of the lightest vector mesons $m_{\text{V}}$ is 2.2 times the mass of the lightest pseudoscalars $m_{\text{PS}}$. We will assume that the DM is made of the lightest pseudoscalars. We will use the freeze-in process to produce DM, mainly following the framework discussed in [24, 25]. It is worth mentioning that if the fermions have light masses in the UV [26, 27], the DM could become the dilaton forbidden dark matter, as discussed in [28].
@@
$\mathcal{L}_{\text{SM}}+\mathcal{L}_{\text{CFT}}+\sum\lambda\frac{\mathcal{O}_{\text{H}}\mathcal{O}_{\Delta}}{\Lambda^{d}},\ d=d_{\text{SM}}+\Delta-4.$ (11)
Here $\mathcal{O}_{\text{SM}}$ is an SM singlet operator with dimension $d_{\text{SM}}$ and we will work in the regime of small dimensionless Wilson coefficient $\lambda\ll 1$. In the Higgs portal, the choice of scalar operators $\mathcal{O}_{\text{SM}}$ in the UV up to dimension four can be $|\text{H}|^{2}$ or $|D_{\mu}\text{H}|^{2}$. From [25] and also later, when $d>1/2$, the freeze-in processes are UV-dominated, and when $d<1/2$, the processes are IR-dominated. We assume the reheating temperature $T_{R}$ to be the highest temperature of the thermal bath, and it is much higher than the electroweak (EW) symmetry breaking scale $\sim 100\text{ GeV}$. The UV-dominating interaction will be $|D_{\mu}\text{H}|^{2}\mathcal{O}_{\psi}$, whose $d=1.8$. A natural choice for the effective coupling is $\Lambda^{\Delta}\equiv\Lambda_{\text{cw}}^{\Delta}$. The interaction $|D_{\mu}\text{H}|^{2}\mathcal{O}_{F}$, and the interactions between currents and stress tensors, e.g. $J_{\text{SM}}J_{\text{DS}}$, $T_{\text{SM}}T_{\text{DS}}$, etc., will be sub-dominant, because of the large scaling dimensions. As for the interaction $|\text{H}|^{2}\mathcal{O}_{F}$, the effective coupling is natural to be $\Lambda^{\Delta-2}\equiv\Lambda_{\text{cw}}^{\Delta}/\langle v\rangle^{2}$, where $\langle v\rangle$ is the vacuum expectation value (VEV) of the Higgs. This coupling is much suppressed in the UV regime. Nonetheless, the indirect interaction between the Higgs and $\mathcal{O}_{F}$ will also be sub-dominate and it is natural not to include this interaction in the UV freeze-in processes.
In the IR-dominant regime, the effective interaction will become
@@
We assume that the Dark Walker completely breaks the conformal symmetries and develops a mass gap $m_{\text{PS}}$ at scale $T_{D}\sim m_{\text{PS}}$. The SM can be treated as a free CFT with $m_{\text{PS}}\lesssim 100\text{ GeV}$ [24]. Thus the UV DS production can be treated as the Dark Walker CFT states from the SM CFT. From the classification of mass generation mechanisms in [14], we know that in the SMG case, the system is anomaly-free, and potentially the effective interactions between the pseudoscalars will be strongly coupled, since there is no small scale to suppress the interactions. The particle number density in the DS will thus not be conserved, and a more relevant quantity is the energy density flowing into the DS. Assuming that the DS evolves relativistically, the energy density $\rho_{\text{CFT}}$ of the DS CFT is controlled by the Boltzmann equation:
@@
By considering the $2\rightarrow\text{CFT}$ scattering process $|D_{\mu}\text{H}|^{2}\rightarrow\text{CFT}$, [24] has computed the energy injection rate coefficient $\kappa_{1}$ to be
@@
The IR-dominant interaction (12) will give the Higgs decay contribution [6, 7, 25]
@@
where $f_{\Delta}\equiv 4^{-\Delta}\pi^{1/2-2\Delta}\Gamma(\Delta+1/2)/(\Gamma(\Delta-1)\Gamma(2\Delta))$ and $K_{\nu}(x)$ is the modified Bessel functions of the second kind. The process roughly starts around the EW scale $\sim v$ and ends when the SM temperature reaches the Higgs mass $m_{h}$ [25]. After the temperature reaches the Higgs mass, the Higgs decouples from the thermal bath and the decay will be exponentially suppressed. We can take the limit $m_{h}\rightarrow 0$ in the thermal average computation and simplify the decay contribution:
@@
We can see that the mass of DM particles increases with larger $\Lambda_{\text{cw}}$ and smaller $T_{R}$. By estimating the DM lifetime $\tau_{\text{DM}}\sim M_{\text{pl}}^{4}/m_{\text{PS}}^{5}>10^{28}\text{ s}$, we can setup the upper bound for the DM mass once we consider the decay into gravitons and $\gamma$-rays [24, 29]. A lower bound can be established by considering standard DM self-interaction, taking $\langle\sigma v\rangle\sim\pi/m_{\text{PS}}^{2}$. The region $m_{\text{PS}}\lesssim 10^{-1}\text{ GeV}$ will be excluded from the Bullet Cluster observations [24]. Assuming that the reheating temperature cannot exceed the inflationary Hubble constant and taking the upper and lower bounds of DM mass, we can plot the region of permitted masses, as shown in Figure 4.

### v4

@@ Abstract
Introduction.—Assuming that the early universe underwent a phase of cosmic inflation, fields that weakly couple to the inflaton can play a crucial role in shaping the primordial non-Gaussianities—also referred to as cosmological correlators. One possible origin of scalar perturbations during inflation is the presence of composite particles, which arise from strongly coupled dynamics among more elementary constituents, similar to the mechanisms in quantum chromodynamics (QCD) and models of the composite Higgs. This scenario has been far less explored compared to the more conventional treatments of cosmological correlators involving fundamental particle exchanges. For recent developments in this direction, see, for example, [Green:2013rd, Baumgart:2021ptt, Aoki:2023tjm, Chakraborty:2025myb, Pimentel:2025rds].
@@
Setup.—To investigate a strongly coupled DS during inflation, it is useful to classify different scenarios according to the mass gap: for example, $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$, $\text{M}_{\text{Gap}}\sim H_{\text{inf}}$ and $\text{M}_{\text{Gap}}\gg H_{\text{inf}}$. In this work, the Dark Walker model falls into the $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$ regime, where the DS is well approximated by a conformal field theory. Such a conformal sector may contain operators with large anomalous dimensions, which are called unparticles, as introduced in [Georgi:2007ek, Grinstein:2008qk], etc. An extended notion of unparticles with a mass gap has also been discussed in [Strassler:2008bv], but lies beyond the scope of this work. Gapless unparticles, however, cannot serve as DM, since their behavior resembles radiation rather than particles. Thus, the Dark Walker has to experience a phase transition, forming gapped states. During inflation, the model remains near the infrared fixed point (IRFP) long enough to leave imprints in the primordial non-Gaussianity. This behavior exemplifies a form of “walking dynamics”, which gives rise to the name “Dark Walker”.
@@
It is long known that there exists a non-trivial IRFP for $\text{SU}(N_{c})$ theories with a certain range of $N_{f}$, which is called the conformal window [Banks:1981nn]. The $\beta$-function of $\text{SU}(N_{c})$ theories in the asymptotically free phase is given in [Gross:1973id, Caswell:1974gg]
@@
From the $\beta$-function we notice that, generally in the perturbative regime $g\lesssim 4\pi$, the IRFP is stable [Banks:1981nn], which means that a small relevant deformation will not drag the Dark Walker into the gapped phase. Therefore, we have to look for special gauge theories. Fortunately, recent studies in lattice QCD (e.g. [Butt:2024kxi, Witzel:2024bly]) have revealed that, for the SU(2) theory with four flavors of massless Dirac fermions and the $\text{SU}(3)$ theory with eight flavors of massless Dirac fermions, the ultraviolet (UV) and IR fixed points will merge, as illustrated in Figure 1.
@@
In the following, we will take the Dark Walker to be an $\text{SU}(3)_{D}$ theory with $N_{f}=8$ flavors of massless Dirac fermions. This is an illustrator of our Dark Walker mechanism. In general, this scenario should hold for all gauge models with a merged fixed point (MFP). This kind of mass generation is called the symmetric mass generation (SMG), a phenomenon where the fermions acquire mass gaps while the chiral symmetry remains unbroken. For a recent review, the readers can refer to [Wang:2022ucy].
During Inflation.—In this work, we choose the value of $H_{\text{inf}}$ to be the current upper bound value from [Planck:2018jri], which is $H_{\text{inf}}\sim 2.5\times 10^{-5}M_{\text{pl}}\sim 6.0\times 10^{13}\text{ GeV}$. In the context of cosmology, it is convenient to use the number of $e$-folds NN to measure the ratio between scales. There are $\sim 10$ $e$-folds between the Planck mass scale and the Hubble scale during inflation. Once we assume the inflation starts from the $M_{\text{pl}}$, the working energy scale during inflation flows from $M_{\text{pl}}$ to $H_{\text{inf}}$, and stays at $H_{\text{inf}}$ until the end of inflation.
@@
We will take the critical gauge coupling $g_{c}^{2}$ to be the lattice result of [Witzel:2024bly]. The critical gradient-flow coupling $g_{\text{GF}}^{2}\approx 25$ will give us $g_{c}^{2}\lesssim 12$, with reference to [Luscher:2010iy]. Surely, this one-loop matching in [Luscher:2010iy] will break up for strong couplings, but all we need is an upper bound for the number of $e$-folds needed. We will restrict the ranges of $N_{\text{cw}}$ and the initial value of the gauge coupling $g_{0}$ after constructing the effective interactions.
Once our model reaches the MFP, the gauge theory (1) becomes a conformal theory, and we cannot use this Lagrangian. Generally, it is not easy to derive the operator spectrum and the operator product expansion (OPE) coefficients between operators from the UV theory (1). Nevertheless, we can focus on a few operators that we expect to see from the UV theory. An important one is the fermion bilinear operator $\mathcal{O}_{\psi}\equiv\bar{\psi}\psi$, as the anomalous dimension $\gamma_{m}$ of $\mathcal{O}_{\psi}$ will be greater than 1 near the MFP, studied in [Cheng:2013eu]. This property will trigger “walking” and drag the Dark Walker into the gapped phase, as we will see below. Another possible primary scalar operator is $\mathcal{O}_{F}\equiv\text{Tr}F^{2}$, where the scaling dimension $\Delta_{F}\lesssim 4$. The explicit value can only be found from lattice simulations. The operator $\mathcal{O}_{\psi}$ has the scaling dimension $\Delta_{\psi}=3-\gamma_{m}\approx 1.8$ from [Cheng:2013eu], and we assume $\Delta_{F}\approx 3.9$ for $\mathcal{O}_{F}$.
Using the language of effective field theory (EFT) of inflation [Cheung:2007st, Lee:2016vti], we can construct the linear interactions between the scalar primary operators $\mathcal{O}_{\Delta}$ and the scalar perturbation $\pi$ as given in [Green:2013rd, Pimentel:2025rds]:
@@
In addition, assuming that $\langle\mathcal{O}_{\Delta}^{3}\rangle$ is negligible, the bispectrum is determined by the quadratic derivative coupling $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\Delta}$, as discussed in [Pimentel:2025rds]. Having established the EFT of inflation with unparticles, we now turn to its cosmological signatures.
Non-Gaussianity.—The cosmological bispectrum can be calculated using the results in [Pimentel:2025rds]. The non-Gaussianity $f_{\text{NL}}$ is defined as
@@
is the three-point correlation function of curvature perturbation $\zeta$. $\Delta_{\zeta}^{2}=2.1\times 10^{-9}$ is the dimensionless power spectrum of $\zeta$ from [Planck:2018vyg]. At linear order, the curvature perturbation $\zeta$ relates to the scalar perturbation $\pi$ via
@@
Taking the slow-roll parameter $\epsilon$ to be $0.0097$ according to [Planck:2018jri], we find that $1.2\times 10^{-8}\lesssim f_{\text{NL},\psi}\lesssim-13$ and $|f_{\text{NL},F}|<10^{-10}$. The full expression of $f_{\text{NL}}$ is presented in Appendix A. The $f_{\text{NL}}$ with $\mathcal{O}_{\psi}$ is desirable because it has the potential to be $O(1)$, which is around the center of the latest observation $f_{\text{NL}}=6\pm 46$ [Jung:2025nss], while large enough to be detectable with the experiments in the near future. The $f_{\text{NL},F}$ is too small to be detected. Even the futuristic 21-cm tomography can only give a limiting strength of non-Gaussianity for observability about $f_{\text{NL}}\gtrsim 10^{-4}$ [Meerburg:2016zdz]. Notice that the choice of $H_{\text{inf}}$ will significantly affect the value of $f_{\text{NL}}$, as indicated in Figure 2.
@@
Figure 3 shows the shape functions with $\mathcal{O}_{\psi}$ and $\mathcal{O}_{F}$ exchanges. Both of them are near-equilateral, while the squeezed limit behaviors degenerate with the equilateral non-Gaussianity. Only the full shapes of the bispectra can break this degeneracy, as emphasized in [Pimentel:2025rds].
@@
After inflation.—This section connects the inflationary phase to the present-day dark matter abundance. After the inflation ends, the universe becomes almost flat. The interactions $\dot{\pi}\mathcal{O}_{\psi}$ and $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\psi}$ will freeze, since the energy scale becomes too small compared to the EFT of inflation scale $M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}$. However, even though the linear mixing interaction $\dot{\pi}\mathcal{O}_{\psi}$ remains small before it freezes, its nature as a relevant operator will drag our model into the gapped phase. The gapped states are natural DM candidates. The masses of the lowest lying states of pseudoscalars (PSs) and vector mesons (Vs) can be read from the lattice QCD results [Witzel:2024bly, Rinaldi:2015axa], where the mass of the lightest vector mesons $m_{\text{V}}$ is 2.2 times the mass of the lightest pseudoscalars $m_{\text{PS}}$. We will assume that the DM is made of the lightest pseudoscalars. We will use the freeze-in process to produce DM, mainly following the framework discussed in [Redi:2021ipn, Hong:2019nwd]. It is worth mentioning that if the fermions have light masses in the UV [LatKMI:2016xxi, LatticeStrongDynamics:2023bqp], the DM could become the dilaton forbidden dark matter, as discussed in [Appelquist:2024koa].
@@
$\mathcal{L}_{\text{SM}}+\mathcal{L}_{\text{CFT}}+\sum\lambda\frac{\mathcal{O}_{\text{SM}}\mathcal{O}_{\Delta}}{\Lambda^{d}},\ d=d_{\text{SM}}+\Delta-4.$ (11)
Here $\mathcal{O}_{\text{SM}}$ is an SM singlet operator with dimension $d_{\text{SM}}$ and we will work in the regime of small dimensionless Wilson coefficient $\lambda\ll 1$. In the Higgs portal, the choice of scalar operators $\mathcal{O}_{\text{SM}}$ in the UV up to dimension four can be $|\text{H}|^{2}$ or $|D_{\mu}\text{H}|^{2}$. From [Hong:2019nwd] and also later, when $d>1/2$, the freeze-in processes are UV-dominated, and when $d<1/2$, the processes are IR-dominated. We assume the reheating temperature $T_{R}$ to be the highest temperature of the thermal bath, and it is much higher than the electroweak (EW) symmetry breaking scale $\sim 100\text{ GeV}$. The UV-dominating interaction will be $|D_{\mu}\text{H}|^{2}\mathcal{O}_{\psi}$, whose $d=1.8$. A natural choice for the effective coupling is $\Lambda^{\Delta}\equiv\Lambda_{\text{cw}}^{\Delta}$. The interaction $|D_{\mu}\text{H}|^{2}\mathcal{O}_{F}$, and the interactions between currents and stress tensors, e.g. $J_{\text{SM}}J_{\text{DS}}$, $T_{\text{SM}}T_{\text{DS}}$, etc., will be sub-dominant, because of the large scaling dimensions. As for the interaction $|\text{H}|^{2}\mathcal{O}_{F}$, the effective coupling is natural to be $\Lambda^{\Delta-2}\equiv\Lambda_{\text{cw}}^{\Delta}/\langle v\rangle^{2}$, where $\langle v\rangle$ is the vacuum expectation value (VEV) of the Higgs. This coupling is much suppressed in the UV regime. Nonetheless, the indirect interaction between the Higgs and $\mathcal{O}_{F}$ will also be sub-dominate and it is natural not to include this interaction in the UV freeze-in processes.
In the IR-dominant regime, we expand Higgs near the VEV as $\text{H}=(v+h)/\sqrt{2}$ in (11). The dominant effective interaction will become
@@
We assume that the Dark Walker completely breaks the conformal symmetries and develops a mass gap $m_{\text{PS}}$ at scale $T_{D}\sim m_{\text{PS}}$. The SM can be treated as a free CFT with $m_{\text{PS}}\lesssim 100\text{ GeV}$ [Redi:2021ipn]. Thus the UV DS production can be treated as the Dark Walker CFT states from the SM CFT. From the classification of mass generation mechanisms in [Wang:2022ucy], we know that in the SMG case, the system is anomaly-free, and potentially the effective interactions between the pseudoscalars will be strongly coupled, since there is no small scale to suppress the interactions. The particle number density in the DS will thus not be conserved, and a more relevant quantity is the energy density flowing into the DS. Assuming that the DS evolves relativistically, the energy density $\rho_{\text{CFT}}$ of the DS CFT is controlled by the Boltzmann equation:
@@
By considering the $2\rightarrow\text{CFT}$ scattering process $|D_{\mu}\text{H}|^{2}\rightarrow\text{CFT}$, [Redi:2021ipn] has computed the energy injection rate coefficient $\kappa_{1}$ to be
@@
The IR-dominant interaction (12) will give the Higgs decay contribution [Georgi:2007ek, Grinstein:2008qk, Hong:2019nwd]
@@
where $f_{\Delta}\equiv 4^{-\Delta}\pi^{1/2-2\Delta}\Gamma(\Delta+1/2)/(\Gamma(\Delta-1)\Gamma(2\Delta))$ and $K_{\nu}(x)$ is the modified Bessel functions of the second kind. The process roughly starts around the EW scale $\sim v$ and ends when the SM temperature reaches the Higgs mass $m_{h}$ [Hong:2019nwd]. After the temperature reaches the Higgs mass, the Higgs decouples from the thermal bath and the decay will be exponentially suppressed. We can take the limit $m_{h}\rightarrow 0$ in the thermal average computation and simplify the decay contribution:
@@
We can see that the mass of DM particles increases with larger $\Lambda_{\text{cw}}$ and smaller $T_{R}$. If we take the lowest value of $\Lambda_{\text{cw}}$, the DM mass becomes
$m_{\text{PS}}\sim 2.5\times 10^{63}\left(\frac{\text{GeV}^{2.6}}{\lambda^{2}T_{R}^{2.6}}\right)^{7/4}\text{ GeV}.$ (26)
By estimating the DM lifetime $\tau_{\text{DM}}\sim M_{\text{pl}}^{4}/m_{\text{PS}}^{5}>10^{28}\text{ s}$, we can setup the upper bound for the DM mass once we consider the decay into gravitons and $\gamma$-rays [Redi:2021ipn, Blanco:2018esa]. The decay into Higgs and other SM particles will be much suppressed by the small portal coupling, as for typical freeze-in processes, and thus will give a looser bound. A lower bound can be established by considering standard DM self-interaction, taking $\langle\sigma v\rangle\sim\pi/m_{\text{PS}}^{2}$. The region $m_{\text{PS}}\lesssim 10^{-1}\text{ GeV}$ will be excluded from the Bullet Cluster observations [Redi:2021ipn]. Assuming that the reheating temperature cannot exceed the inflationary Hubble constant and taking the upper and lower bounds of DM mass, we can plot the region of permitted masses, as shown in Figure 4.